# Quantum Algebraic Diversity: Single-Copy Density Matrix Estimation via Group-Structured Measurements

### 04 Apr 2026 v1

@@ Abstract
We extend the algebraic diversity (AD) framework from classical signal processing to quantum measurement theory. The central result—the Quantum Algebraic Diversity (QAD) Theorem—establishes that a group-structured positive operator-valued measure (POVM) applied to a single copy of a quantum state produces a group-averaged density matrix estimator that recovers the spectral structure of the true density matrix, analogous to the classical result that a group-averaged outer product recovers covariance eigenstructure from a single observation. We establish a formal Classical-Quantum Duality Map connecting classical covariance estimation to quantum state tomography, and prove an Optimality Inheritance Theorem showing that classical group optimality transfers to quantum settings via the Born map. SIC-POVMs are identified as algebraic diversity with the Heisenberg-Weyl group, and mutually unbiased bases (MUBs) as algebraic diversity with the Clifford group, revealing the hierarchy $\mathrm{HW}(d)\subseteq\mathcal{C}(d)\subseteq S_{d}$ that mirrors the classical hierarchy $\mathbb{Z}_{M}\subseteq G_{\min}\subseteq S_{M}$. The double-commutator eigenvalue theorem provides polynomial-time adaptive POVM selection. A worked qubit example demonstrates that the group-averaged estimator from a single Pauli measurement recovers a full-rank approximation to a mixed qubit state, achieving fidelity 0.91 where standard single-basis tomography produces a rank-1 estimate with fidelity 0.71. Monte Carlo simulations on qudits of dimension $d=2$ through $d=13$ (200 random states per dimension) confirm that the Heisenberg-Weyl QAD estimator maintains fidelity above 0.90 across all dimensions from a single measurement outcome, while standard tomography fidelity degrades as $\sim 1/d$, with the improvement ratio scaling linearly with $d$ as predicted by the $O(d)$ copy reduction theorem.
@@ I. Introduction
Quantum state tomography—reconstructing the density matrix $\rho$ of a quantum system from measurement outcomes—is a fundamental task in quantum information science. For a $d$-dimensional system, the density matrix has $d^{2}-1$ independent real parameters. Standard tomography requires $O(d^{2}/\varepsilon^{2})$ identically prepared copies for trace-distance error $\varepsilon$ [3], a cost that scales prohibitively for multi-qubit systems.
In classical signal processing, an analogous bottleneck exists: estimating the $M\times M$ covariance matrix of a signal requires $L\geq 2M$ independent observations. The algebraic diversity framework [1] dissolves this bottleneck by proving that the covariance eigenstructure can be recovered from a single observation via a group-averaged outer product. The key mechanism is that each group element provides an algebraically distinct “view” of the observation, and averaging over the group orbit separates the structured (signal) component from the unstructured (noise) component.
This paper shows that the same algebraic principle extends to quantum measurement. The measurement outcome distribution—the Born probability vector—plays the role of the classical observation, and the density matrix plays the role of the covariance matrix. A group-structured POVM applied to a single state copy produces a group-averaged estimator that recovers the spectral structure of $\rho$, just as the classical group-averaged outer product recovers the spectral structure of $\mathbf{R}$.
@@ III. The Classical-Quantum Duality Map
The classical group-averaged estimator (1) is a function of the group action and the outer product. Under the duality map, the outer product $\mathbf{x}\mathbf{x}^{*}$ becomes the projector $|\phi_{m}\rangle\langle\phi_{m}|$ (the rank-1 operator associated with the measurement outcome), and the group action $\pi_{g}$ becomes the unitary rotation $U_{g}(\cdot)U_{g}^{\dagger}$. The algebraic structure—group averaging of rank-1 operators—is identical.
@@ IV. The QAD Theorem
    Copy reduction: The group-averaged estimator from $N_{\mathrm{AD}}$ copies achieves trace-distance error $\varepsilon$ when $N_{\mathrm{AD}}=O(d/\varepsilon^{2})$, compared to $N_{\mathrm{std}}=O(d^{2}/\varepsilon^{2})$ for standard tomography—a factor of $d$ reduction.
@@
Part (iii). The classical processing gain of algebraic diversity is $10\log_{10}(M)$ dB from $M$ group elements [1]. In the quantum setting, $|G|\geq d$ group elements produce a processing gain of $10\log_{10}(d)$ dB per copy. The trace-distance error of the group-averaged estimator from NN copies scales as $\varepsilon\sim 1/\sqrt{Nd}$ (the factor of $d$ arising from the processing gain), compared to $\varepsilon\sim 1/\sqrt{N}$ for standard tomography. Setting $\varepsilon=1/\sqrt{N_{\mathrm{AD}}\cdot d}=1/\sqrt{N_{\mathrm{std}}}$ gives $N_{\mathrm{std}}=d\cdot N_{\mathrm{AD}}$. ∎
@@ VI. Optimality Inheritance
By the Commutativity–KL Equivalence [1], the group minimizing $\delta$ produces the spectral decomposition closest to the KL (optimal) decomposition. Since the KL optimality properties—variance concentration, orthogonality, minimum reconstruction error—depend only on the algebraic relationship between the group and the matrix, they hold identically for $(\pi_{g},\mathbf{R})$ and $(U_{g},\rho)$ when the two pairs have isomorphic algebraic structure.
@@ C. QAD with Pauli Group
This is the maximally mixed state—full rank but uninformative. This reflects the fact that $\{I,X,Y,Z\}$ acting on $|0\rangle$ generates the orbit $\{|0\rangle,|1\rangle,i|1\rangle,|0\rangle\}$, which spans $\mathbb{C}^{2}$ but symmetrizes away all structure because the Pauli group is too large relative to the state’s symmetry.
@@ F. Summary
The example illustrates the core principle: the group must be (a) large enough to generate a spanning orbit (producing a full-rank estimator) and (b) small enough that its structure is matched to $\rho$ (preserving spectral information). The Hadamard group $\{I,H\}$ achieves both, producing a full-rank estimate with fidelity 0.91 from a single measurement outcome—a substantial improvement over the rank-1, fidelity-0.80 estimate of standard single-basis tomography.
@@ A. Setup
    QAD with matched cyclic group: The group-averaged estimator with $G=\mathbb{Z}_{d}$ conjugated into the eigenbasis of $\rho$ (oracle benchmark—requires knowledge of $\rho$’s eigenvectors).
@@ B. Results
Table 1 and Fig. 1 present the results.
@@
Figure 1: Quantum algebraic diversity on qudits. (a) Mean fidelity from a single measurement outcome: HW($d$) QAD maintains $F>0.90$ across all dimensions while standard single-basis fidelity collapses as $\sim 1/d$. (b) Fidelity improvement ratio scales linearly with $d$, confirming the $O(d)$ copy reduction of Theorem 6(iii). (c) Spectral recovery error (eigenvalue $\ell_{2}$ distance) for HW and matched groups. 200 random mixed states per dimension, purity $\approx 0.7$.
@@
Finding 1: QAD fidelity is dimension-independent. The HW($d$) group-averaged estimator maintains mean fidelity $F>0.90$ from $d=2$ through $d=13$, while the standard rank-1 projector fidelity degrades as approximately $1/d$ (from 0.514 at $d=2$ to 0.079 at $d=13$). This confirms the full-rank property of Theorem 6(i): the $d^{2}$ group elements of $\mathrm{HW}(d)$ generate a spanning orbit that produces a full-rank estimator regardless of dimension.
Finding 2: The improvement ratio scales linearly with $d$. The fidelity ratio HW/Standard grows from $1.9\times$ at $d=2$ to $11.3\times$ at $d=13$, consistent with the $O(d)$ copy reduction predicted by Theorem 6(iii). Extrapolating, a 10-qubit system ($d=1024$) would achieve approximately $1000\times$ improvement—reducing the copy requirement from $O(d^{2})\approx 10^{6}$ to $O(d)\approx 10^{3}$.
Finding 3: HW outperforms the matched cyclic group. The Heisenberg-Weyl group ($|G|=d^{2}$) consistently outperforms the matched cyclic group ($|G|=d$) because the larger group provides more algebraic views of the measurement outcome. This is consistent with the classical result that larger matched groups yield better estimation, up to the universal optimum of $S_{d}$. The matched cyclic group, despite using oracle knowledge of $\rho$’s eigenbasis, achieves lower fidelity because its order-$d$ orbit provides only $d$ distinct views, compared to $d^{2}$ views from HW($d$).
@@ IX. Implications for Quantum Error Correction
Quantum error correction codes are defined by stabilizer groups—subgroups of the Pauli group that fix the code space. In the QAD framework, a stabilizer code is a quantum state whose density matrix commutes with the stabilizer group: $S,\rho_{\mathrm{code}}]=0$ for all stabilizers $S$.
The QAD Theorem implies that syndrome extraction—measuring which stabilizer is violated—is an instance of algebraic diversity with the stabilizer group as the measurement group. The group-averaged estimator with the stabilizer group achieves optimal syndrome extraction from minimal copies, because the commutativity condition is exactly satisfied (the stabilizer group is, by definition, the matched group for the code space).
@@
X. Discussion

### 13 Apr 2026 v2

@@ Abstract
We extend the algebraic diversity (AD) framework from classical signal processing to quantum measurement theory. The central result, the Quantum Algebraic Diversity (QAD) Theorem, establishes that a group-structured positive operator-valued measure (POVM) applied to a single copy of a quantum state produces a group-averaged density matrix estimator that recovers the spectral structure of the true density matrix, analogous to the classical result that a group-averaged outer product recovers covariance eigenstructure from a single observation. We establish a formal Classical-Quantum Duality Map connecting classical covariance estimation to quantum state tomography, and prove an Optimality Inheritance Theorem showing that classical group optimality transfers to quantum settings via the Born map. SIC-POVMs are identified as algebraic diversity with the Heisenberg-Weyl group, and mutually unbiased bases (MUBs) as algebraic diversity with the Clifford group, revealing the hierarchy $\mathrm{HW}(d)\subseteq\mathcal{C}(d)\subseteq S_{d}$ that mirrors the classical hierarchy $\mathbb{Z}_{M}\subseteq G_{\min}\subseteq S_{M}$. The double-commutator eigenvalue theorem provides polynomial-time adaptive POVM selection. A worked qubit example demonstrates that the group-averaged estimator from a single Pauli measurement recovers a full-rank approximation to a mixed qubit state, achieving fidelity 0.91 where standard single-basis tomography produces a rank-1 estimate with fidelity 0.71. Monte Carlo simulations on qudits of dimension $d=2$ through $d=13$ (200 random states per dimension) confirm that the Heisenberg-Weyl QAD estimator maintains fidelity above 0.90 across all dimensions from a single measurement outcome, while standard tomography fidelity degrades as $\sim 1/d$, with the improvement ratio scaling linearly with $d$ as predicted by the $O(d)$ copy reduction theorem. We define the structural capacity of a density matrix $\kappa(\rho)=1+1/\operatorname{Tr}(\rho^{2})$, governed by the Renyi-2 entropy of the eigenvalue distribution, and establish a Structural Measurement Bound: the maximum exploitable structure from a single copy is $\kappa(\rho)$. This parallels Holevo’s bound on extractable content (governed by von Neumann entropy, Renyi order 1) with a bound on exploitable structure (governed by Renyi order 2), revealing that QAD is the estimation-theoretic completion of von Neumann’s measurement theory.
@@ I. Introduction
Quantum state tomography, the reconstruction of the density matrix $\rho$ of a quantum system from measurement outcomes, is a fundamental task in quantum information science. For a $d$-dimensional system, the density matrix has $d^{2}-1$ independent real parameters. Standard tomography requires $O(d^{2}/\varepsilon^{2})$ identically prepared copies for trace-distance error $\varepsilon$ [4], a cost that scales prohibitively for multi-qubit systems.
In classical signal processing, an analogous bottleneck exists: estimating the $M\times M$ covariance matrix of a signal requires $L\geq 2M$ independent observations. The algebraic diversity framework [1] reveals that this requirement is an artifact of using the trivial group $G=\{e\}$ when forming the outer product $\mathbf{x}\mathbf{x}^{H}$. By selecting a richer group of order $M$, the covariance eigenstructure can be recovered from a single observation via a group-averaged outer product. The key mechanism is that each group element provides an algebraically distinct “view” of the observation, and averaging over the group orbit separates the structured (signal) component from the unstructured (noise) component.
This paper shows that the same algebraic principle extends to quantum measurement. The measurement outcome distribution (the Born probability vector) plays the role of the classical observation, and the density matrix plays the role of the covariance matrix. A group-structured POVM applied to a single state copy produces a group-averaged estimator that recovers the spectral structure of $\rho$, just as the classical group-averaged outer product recovers the spectral structure of $\mathbf{R}$.
Quantum algebraic diversity is not an alternative to standard tomography; it is a generalization. To see why, consider three scenarios for estimating the density matrix of a $d$-dimensional quantum system.
In standard tomography, a single copy is measured in a fixed basis $\{|k\rangle\}$. The trivial group $G=\{e\}$ acts on the measurement: no basis rotation is applied. The outcome is a single projector $|k\rangle\langle k|$, which has rank one and contains no information about the off-diagonal structure of $\rho$. To recover $\rho$, one must prepare $O(d^{2})$ identical copies, measure each in a (possibly different) fixed basis, and accumulate the rank-1 outcomes. This is the quantum analog of temporal averaging with the trivial group.
Now suppose the experimenter has access to a SIC-POVM, a symmetric informationally complete measurement with $d^{2}$ outcomes. A SIC-POVM is generated by applying the $d^{2}$ elements of the Heisenberg-Weyl group $\mathrm{HW}(d)$ to a single fiducial vector. The group-averaged estimator from a single measurement outcome has rank $d$ and maintains fidelity above 0.90 across all dimensions tested (Table 1, Figs. 13).
The key observation is that the improvement arises from the same mechanism as spatial beamforming in classical signal processing. The $d^{2}$ group elements of $\mathrm{HW}(d)$ provide $d^{2}$ algebraically distinct views of the single measurement outcome, just as $M$ sensors provide $M$ independent views of a signal. The structured component (the density matrix eigenstructure) combines coherently across these views while the measurement noise averages down. Standard tomography uses the trivial group and therefore requires $O(d^{2})$ physical copies to achieve what $\mathrm{HW}(d)$ achieves algebraically from one copy. The copies do not provide new physics; they compensate for the absence of group structure in the measurement.
@@ A. Contributions
  6. 6.
    Structural Capacity and von Neumann–Holevo Connection: We define the structural capacity $\kappa(\rho)=1+1/\operatorname{Tr}(\rho^{2})$ of a density matrix and establish a Structural Measurement Bound paralleling Holevo’s bound: von Neumann entropy ($H_{1}$) limits extractable content, structural capacity ($H_{2}$) limits exploitable structure.
@@ II. Classical Algebraic Diversity (Review)
  5. (R5)
    General Algebraic Averaging [2]: The algebraic diversity principle extends beyond the outer product to arbitrary $G$-compatible statistics. The variance reduction is governed by the effective dimension $d_{\mathrm{eff}}(G,f)=|\{f(\pi_{g}(\mathbf{x})):g\in G\}|$ (the number of distinct values the statistic produces on the group orbit). For the outer product, $d_{\mathrm{eff}}=|G|$; the density matrix, as a rank-2 tensor, admits $d_{\mathrm{eff}}=d^{2}$ under the Heisenberg-Weyl group. The classical law of large numbers is recovered as the trivial-group case $d_{\mathrm{eff}}=1$.
@@ III. The Classical-Quantum Duality Map
The classical group-averaged estimator (1) is a function of the group action and the outer product. Under the duality map, the outer product $\mathbf{x}\mathbf{x}^{*}$ becomes the projector $|\phi_{m}\rangle\langle\phi_{m}|$ (the rank-1 operator associated with the measurement outcome), and the group action $\pi_{g}$ becomes the unitary rotation $U_{g}(\cdot)U_{g}^{\dagger}$. The algebraic structure, group averaging of rank-1 operators, is identical.
@@ IV. The QAD Theorem
    Copy reduction: The group-averaged estimator from $N_{\mathrm{AD}}$ copies achieves trace-distance error $\varepsilon$ when $N_{\mathrm{AD}}=O(d/\varepsilon^{2})$, compared to $N_{\mathrm{std}}=O(d^{2}/\varepsilon^{2})$ for standard tomography, a factor of $d$ reduction.
@@
Part (iii). The classical group gain of algebraic diversity is $10\log_{10}(M)$ dB from $M$ group elements [1]. In the quantum setting, $|G|\geq d$ group elements produce a group gain of $10\log_{10}(d)$ dB per copy. The trace-distance error of the group-averaged estimator from $N$ copies scales as $\varepsilon\sim 1/\sqrt{Nd}$ (the factor of $d$ arising from the group gain), compared to $\varepsilon\sim 1/\sqrt{N}$ for standard tomography. Setting $\varepsilon=1/\sqrt{N_{\mathrm{AD}}\cdot d}=1/\sqrt{N_{\mathrm{std}}}$ gives $N_{\mathrm{std}}=d\cdot N_{\mathrm{AD}}$. ∎
@@ VI. Optimality Inheritance
By the Commutativity–KL Equivalence [1], the group minimizing $\delta$ produces the spectral decomposition closest to the KL (optimal) decomposition. Since the KL optimality properties (variance concentration, orthogonality, minimum reconstruction error) depend only on the algebraic relationship between the group and the matrix, they hold identically for $(\pi_{g},\mathbf{R})$ and $(U_{g},\rho)$ when the two pairs have isomorphic algebraic structure.
@@ C. QAD with Pauli Group
This is the maximally mixed state, full rank but uninformative. This reflects the fact that $\{I,X,Y,Z\}$ acting on $|0\rangle$ generates the orbit $\{|0\rangle,|1\rangle,i|1\rangle,|0\rangle\}$, which spans $\mathbb{C}^{2}$ but symmetrizes away all structure because the Pauli group is too large relative to the state’s symmetry.
@@ F. Summary
The example illustrates the core principle: the group must be (a) large enough to generate a spanning orbit (producing a full-rank estimator) and (b) small enough that its structure is matched to $\rho$ (preserving spectral information). The Hadamard group $\{I,H\}$ achieves both, producing a full-rank estimate with fidelity 0.91 from a single measurement outcome, a substantial improvement over the rank-1, fidelity-0.80 estimate of standard single-basis tomography.
@@ A. Setup
    QAD with matched cyclic group: The group-averaged estimator with $G=\mathbb{Z}_{d}$ conjugated into the eigenbasis of $\rho$ (oracle benchmark, requires knowledge of $\rho$’s eigenvectors).
@@ B. Results
Table 1 and Figs. 13 present the results.
@@
Figure 1: Single-copy fidelity vs Hilbert space dimension. The HW($d$) QAD estimator maintains mean fidelity $F>0.90$ from $d=2$ through $d=13$, while standard single-basis fidelity collapses as $\sim 1/d$. The matched cyclic group (using oracle knowledge of $\rho$’s eigenbasis) achieves intermediate fidelity. 200 random mixed states per dimension, purity $\approx 0.7$.
Figure 2: Fidelity improvement ratio over standard tomography. The HW(dd)/Standard ratio grows from $1.9\times$ at $d=2$ to $11.3\times$ at $d=13$, scaling linearly with $d$ and confirming the $O(d)$ copy reduction of Theorem 6(iii). Extrapolating, a 10-qubit system ($d=1024$) would achieve approximately $1000\times$ improvement.
[…]
@@
Finding 1: QAD fidelity is dimension-independent (Fig. 1). The HW($d$) group-averaged estimator maintains mean fidelity $F>0.90$ from $d=2$ through $d=13$, while the standard rank-1 projector fidelity degrades as approximately $1/d$ (from 0.514 at $d=2$ to 0.079 at $d=13$). This confirms the full-rank property of Theorem 6(i): the $d^{2}$ group elements of $\mathrm{HW}(d)$ generate a spanning orbit that produces a full-rank estimator regardless of dimension.
Finding 2: The improvement ratio scales linearly with $d$ (Fig. 2). The fidelity ratio HW/Standard grows from $1.9\times$ at $d=2$ to $11.3\times$ at $d=13$, consistent with the $O(d)$ copy reduction predicted by Theorem 6(iii). Extrapolating, a 10-qubit system ($d=1024$) would achieve approximately $1000\times$ improvement, reducing the copy requirement from $O(d^{2})\approx 10^{6}$ to $O(d)\approx 10^{3}$.
Finding 3: HW outperforms the matched cyclic group (Fig. 3). The Heisenberg-Weyl group ($|G|=d^{2}$) consistently outperforms the matched cyclic group ($|G|=d$) because the larger group provides more algebraic views of the measurement outcome. This is consistent with the classical result that larger matched groups yield better estimation, up to the universal optimum of $S_{d}$. The matched cyclic group, despite using oracle knowledge of $\rho$’s eigenbasis, achieves lower fidelity because its order-$d$ orbit provides only $d$ distinct views, compared to $d^{2}$ views from HW($d$).
@@ IX. Implications for Quantum Error Correction
Quantum error correction codes are defined by stabilizer groups, which are subgroups of the Pauli group that fix the code space. In the QAD framework, a stabilizer code is a quantum state whose density matrix commutes with the stabilizer group: $S,\rho_{\mathrm{code}}]=0$ for all stabilizers $S$.
The QAD Theorem implies that syndrome extraction (measuring which stabilizer is violated) is an instance of algebraic diversity with the stabilizer group as the measurement group. The group-averaged estimator with the stabilizer group achieves optimal syndrome extraction from minimal copies, because the commutativity condition is exactly satisfied (the stabilizer group is, by definition, the matched group for the code space).
@@
X. Structural Capacity and the von Neumann–Holevo Connection
The QAD framework reveals a structural complement to the information-theoretic quantities introduced by von Neumann [9] and bounded by Holevo [10]. The connection rests on the distinction between two entropic functionals of the density matrix eigenvalues, each governing a different aspect of quantum information.
A. Von Neumann Entropy as Content; Structural Capacity as Structure
The von Neumann entropy $S(\rho)=-\operatorname{Tr}(\rho\log\rho)=-\sum_{k}\lambda_{k}\log\lambda_{k}$ is the Shannon entropy ($H_{1}$, the Renyi entropy of order 1) applied to the eigenvalue distribution of $\rho$. It quantifies the information content of the quantum state: the maximum number of classical bits extractable per copy in the asymptotic limit.
The structural capacity of a density matrix, defined in analogy with the classical structural capacity [1], is
$\kappa(\rho)=1+\frac{(\operatorname{Tr}\,\rho)^{2}}{\operatorname{Tr}(\rho^{2})}=1+\frac{1}{\operatorname{Tr}(\rho^{2})},$ (14)
since $\operatorname{Tr}\,\rho=1$. The quantity $1/\operatorname{Tr}(\rho^{2})=\exp(H_{2}(\rho))$ is the exponential of the Renyi-2 entropy of $\rho$’s eigenvalues, and is also the participation ratio of the eigenvalue distribution. The structural capacity quantifies the effective dimensionality of the state: the number of eigenvalues that contribute meaningfully to the spectral structure (Fig. 5). For a pure state, $\operatorname{Tr}(\rho^{2})=1$ and $\kappa=2$ (minimal structure: one active eigenvalue). For the maximally mixed state $\rho=\mathbb{I}_{d}/d$, $\operatorname{Tr}(\rho^{2})=1/d$ and $\kappa=1+d$ (maximal structure: all eigenvalues active).
The two quantities are governed by different Renyi orders of the same eigenvalue distribution (Fig. 4):
Quantity Renyi order Measures
$S(\rho)$ (von Neumann) $\alpha=1$ Information content
$\kappa(\rho)$ (structural capacity) $\alpha=2$ Exploitable structure
This is the quantum manifestation of the classical content-vs-structure duality [1]: Shannon’s theory (order 1) characterizes how much information a source produces; AD’s structural theory (order 2) characterizes how much structure a single observation reveals.
Figure 4: Content vs structure for 200 random mixed states per dimension ($d=2,4,8,13$). Von Neumann entropy $S(\rho)$ (Renyi-1, horizontal) measures information content; structural capacity $\kappa(\rho)$ (Renyi-2, vertical) measures exploitable structure. The two quantities are correlated but not identical: states with equal content can have different structure, and vice versa. Stars: pure states ($S=0$, $\kappa=2$). Large markers with black edges: maximally mixed states ($S=\log_{2}d$, $\kappa=1+d$). Higher-dimensional systems access a larger range of both quantities.
B. The Holevo Bound and the Structural Holevo Bound
Holevo’s theorem [10] establishes the maximum classical information extractable from a quantum state ensemble. For an ensemble $\{p_{x},\rho_{x}\}$ with average state $\bar{\rho}=\sum_{x}p_{x}\rho_{x}$, the accessible information is bounded by the Holevo quantity $\chi=S(\bar{\rho})-\sum_{x}p_{x}S(\rho_{x})$. This is a content bound: it limits how many classical bits can be retrieved per quantum state, using von Neumann entropy ($H_{1}$).
The QAD framework suggests a parallel structural bound. From the Converse Theorem in the classical setting [1], the structural capacity $\kappa$ determines the maximum number of effective spectral parameters recoverable from a single observation. Applying this to the density matrix:
Proposition 12 (Structural Measurement Bound).
The maximum number of effective structural parameters of $\rho$ recoverable from a single-copy measurement is bounded by $\kappa(\rho)=1+1/\operatorname{Tr}(\rho^{2})$. This bound is achieved by measuring in the eigenbasis of $\rho$ (the matched basis), which is the QAD estimator with the optimal group.
The parallel is:
Holevo bound Structural bound
Limits Extractable content Exploitable structure
Functional $S(\rho)$ (von Neumann) $\kappa(\rho)=1+e^{H_{2}(\rho)}$
Renyi order $\alpha=1$ $\alpha=2$
Achieved by HSW coding Matched-group POVM
Just as Holevo’s bound limits content extraction regardless of the measurement strategy, the structural bound limits structure extraction regardless of the group chosen. A pure state has $\kappa=2$ (minimal recoverable structure) and $S=0$ (no extractable content). The maximally mixed state has $\kappa=1+d$ (maximal recoverable structure) and $S=\log d$ (maximal extractable content). The two bounds are complementary: a state can have high content but low structure (a nearly pure state with many off-diagonal coherences), or high structure but low content (a highly mixed state with a rich eigenvalue spectrum but little surprise in each measurement).
Figure 5: Structural capacity $\kappa(\rho)=1+1/\operatorname{Tr}(\rho^{2})$ versus purity for random mixed states across four Hilbert space dimensions. The black curve is the universal relationship $\kappa=1+1/\text{purity}$. All states from all dimensions lie exactly on this curve, confirming that $\kappa$ is determined solely by purity (equivalently, Renyi-2 entropy). Pure states ($\operatorname{Tr}(\rho^{2})=1$) have $\kappa=2$ regardless of $d$. Maximally mixed states ($\operatorname{Tr}(\rho^{2})=1/d$, large markers with black edges) have $\kappa=1+d$. Higher-dimensional systems access lower purity (more mixed states) and therefore higher structural capacity.
C. Non-Orthogonal States and the Measurement Basis Problem
Von Neumann established that non-orthogonal quantum states permit richer encodings than orthogonal states, but Holevo’s theorem limits how much of that richness is retrievable. In the QAD framework, the measurement basis selection problem, which basis should one measure in to learn the most about $\rho$ from a single copy, is precisely the quantum group matching problem. Non-orthogonal measurement bases (such as SIC-POVMs, which consist of $d^{2}$ non-orthogonal vectors) provide more algebraic views of the state than any orthogonal basis ($d$ vectors), at the cost of statistical overlap between outcomes. The Heisenberg-Weyl group generates the maximal number of non-orthogonal views ($d^{2}$) from a single fiducial, which is why SIC-POVMs achieve the highest single-copy fidelity in the QAD simulations.
The set of mutually unbiased bases is the quantum group catalog: each MUB is a “group” for probing $\rho$, and the question “which MUB reveals the most structure?” is $G^{*}=\arg\min_{G}\delta_{Q}(G,\rho)$ in quantum language. This connects QAD to the active research area of optimal quantum state estimation [4], providing a group-theoretic criterion for adaptive POVM selection that complements the information-theoretic criteria in the existing literature.
XI. Discussion